OpenStudy (anonymous):

y=1/2x-3; (-4,8) write the equation of the line which is parallel to the given line and passes through the given point

OpenStudy (anonymous):

write an equation that has a slope of 1/2 and has the points -4 and 8. in point slope form

OpenStudy (anonymous):

medal

OpenStudy (anonymous):

y=1/2x-3 (-4,8) y=mx+b substitute -4 for x and 8 for y and keep the slope b/c you're looking for something parallel to it 8=1/2(-4)+b 8=-2+b 10=b y=1/2x+10 <-- this is the answer